# Asynchronous Predictive Counterfactual Regret Minimization+ Algorithm in Solving Extensive-Form Games

### v1

@@ 5 Experiments
We now evaluate the empirical convergence rates of our algorithms, APCFR+ and SAPCFR+. Additional experiments on (i) the dynamics of $\alpha^{t}_{I}$ in APCFR+, and (ii) the dynamics of these terms: $\sum_{t=1}^{T}\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2},\ \sum_{t=1}^{T}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}$, and $\sum_{t=1}^{T}(\frac{\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2})$ for APCFR+, SAPCFR+, and PCFR+, are in Appendix D. The results in Appendix D show that, for these terms, their values are smaller in our algorithms compared to PCFR+.
Configurations. Our evaluation includes a comparison with several existing PCFR+ variants as well as other classical CFR algorithms. Specifically, we compare APCFR+ and SAPCFR+ against PCFR+ (Farina et al., 2021), Stable PCFR+ (Farina et al., 2023), Smooth PCFR+ (Farina et al., 2023), vanilla CFR (Zinkevich et al., 2007), CFR+ (Tammelin, 2014), and DCFR (Brown & Sandholm, 2019a). Each algorithm is run for 5000 iterations, which is sufficient to achieve low exploitability in the tested games. As done in the original PCFR+ version, we employ alternating updates and quadratic averaging for APCFR+, SAPCFR+, Stable PCFR+, and Smooth PCFR+. In addition, for Stable PCFR+ and Smooth PCFR+, we set their step-sizes to 1, the configuration of their original versions. All experiments are conducted on a machine with a Xeon(R) Gold 6444Y CPU and 256 GB of memory.
Figure 4: Empirical convergence for CFR, CFR+, DCFR, PCFR+, Stable PCFR+, Smooth PCFR+, APCFR+, and SAPCFR+ in HUNL Subgames. Notably, the final exploitability of our algorithms in both subgames is approximately $70\%$ of that of PCFR+.
@@
Empirical convergence rates in standard IIG benchmarks. We now present the empirical convergence rates across five standard IIG benchmarks, namely, Kuhn Poker, Leduc Poker, Goofspiel Poker, Liar’s Dice, and Battleship. All the games are implemented using OpenSpiel (Lanctot et al., 2019). The algorithm implementations are based on the open-source code of LiteEFG (Liu et al., 2024), as LiteEFG provides a significant speedup—approximately 100 times faster in executing the same number of iterations compared to the default implementation in OpenSpiel. The experimental results are shown in Figure 3. We observe that, for most of the tested games, except for Battleship ($3$) and Goofspiel ($4$), our algorithms, APCFR+ and SAPCFR+, significantly outperform PCFR+. Even in Battleship ($3$) and Goofspiel ($4$), APCFR+ and SAPCFR+ demonstrate performance similar to that of PCFR+, achieving the similar exploitability after 5000 iterations. In fact, based on the experimental results in Appendix D (Figures 8 and 9), we observe that the games, such as Battleship (3) and Goofspiel (4), where our algorithms do not show an improvement over PCFR+ are also games where PCFR+ exhibits a rapid decrease in the inaccuracy between the predicted and observed instantaneous counterfactual regrets (detailed discussions are provided in Appendix D). Furthermore, the performance gap between APCFR+ and SAPCFR+ is relatively small. Specifically, APCFR+ is only slightly outperforms SAPCFR+ in Leduc Poker, Battleship (3), Liar’s Dice (5), and Liar’s Dice (6). This small performance gap means that in practical applications, SAPCFR+ can be directly used due to its ease of implementation and a faster convergence rate compared to PCFR+. Regarding other variants of PCFR+, namely Stable PCFR+ and Smooth PCFR+, we find that their empirical convergence performance significantly lags behind that of PCFR+. Additionally, compared to other classical CFR algorithms, our algorithms significantly outperform vanilla CFR in all games. Moreover, compared CFR+ and DCFR, our algorithms significantly outperform them in all games except Leduc Poker and Liar’s Dice ($6$). In fact, with an increased number of iterations, our algorithms would likely surpass these classical CFR algorithms even in Leduc Poker.
Empirical convergence rates in HUNL Subgames. We evaluate the empirical convergence rate in HUNL Subgames, which are significantly larger than the standard IIG benchmarks tested (see Appendix D for details). Since OpenSpiel does not support HUNL Subgames, we utilize Poker RL (Steinberger, 2019). More precisely, our code are based on the code from (Xu et al., 2024b). The code in (Xu et al., 2024b) supports only Subgame 3 and Subgame 4, so we conduct experiments solely on these two HUNL Subgames. The results are shown in Figure 4. As observed in the standard IIG benchmark experiments, our algorithms, APCFR+ and SAPCFR+, outperform PCFR+ in both subgames. In fact, our algorithms exceed all others, except DCFR, in both subgames. Although our algorithms do not outperform DCFR, it is worth noting that no other algorithm surpasses DCFR, whose parameters are specifically fine-tuned for HUNL Subgames. Our method and DCFR are not mutually exclusive and can be combined. Specifically, the core idea of our algorithm—the asynchronization of step-sizes between the updates of implicit and explicit accumulated counterfactual regrets—can be integrated with DCFR’s approach, which discounts prior iterations when calculating accumulated regrets. Consequently, we apply the asynchronization mechanism for step-sizes to enhance PDCFR+ (Xu et al., 2024b), a combination of DCFR and PCFR+, resulting in APDCFR+. Notably, APDCFR+ employs a novel aggressive discount weighting scheme for computing accumulated regrets, differing from the scheme used in the original PDCFR+. The experimental results are presented in Figure 5. APDCFR+ significantly outperforms both PDCFR+ and APCFR+, and even surpasses DCFR in Subgame 3. Then, we combine APDCFR+ with Stable PCFR+ to form Stable APDCFR+, further improving performance, and surpassing DCFR in both HUNL Subgames. We also test the combinations of Stable PCFR+ with PDCFR+ and APCFR+, denoted as Stable PDCFR+ and Stable APCFR+, respectively. However, Stable PDCFR+ and Stable APCFR+ even will fall short of PDCFR+ and APCFR+, respectively, let alone Stable APDCFR+, indicating that each component in Stable APDCFR+ is crucial. We also evaluate Stable SAPDCFR+, the combination of Stable PCFR+, PDCFR+, and SAPCFR+. We observe that Stable SAPDCFR+ underperforms Stable APDCFR+ in both subgames. Thus, we present only the Stable APDCFR+. Details of APDCFR+ and Stable APDCFR+ are in Appendix C.
Figure 5: Empirical convergence for DCFR, PCFR+, PDCFR+, APCFR+, PDCFR+, Stable APCFR+, Stable PDCFR+, APDCFR+, and Stable APDCFR+ in HUNL Subgames.
Table 1: Comparison of running times (in minutes).
PCFR+ APCFR+ SAPCFR+
kuhn Poker 0.0020 0.0033 0.0021
Leduc Poker 0.4252 0.6423 0.4408
Battleship (3) 50.1909 58.4390 50.8602
Liar’s Dice (4) 0.4327 0.5690 0.4308
Liar’s Dice (5) 3.8178 4.0365 3.7922
Liar’s Dice (6) 22.8226 24.7680 22.2554
Goofspiel (4) 0.0290 0.0393 0.0279
Goofspiel (5) 0.9323 1.9760 0.9430
Goofspiel (6) 52.4023 56.2751 51.7243
Subgame3 278.7428 297.3784 276.9873
Subgame4 245.1753 254.4792 249.0042
@@
Running times. We now compare the running time of our algorithms with that of PCFR+, both executed with the same number of iterations (i.e., 5000). The experimental results are shown in Table 1. We observe that the running time of APCFR+ is higher compared to PCFR+, primarily due to the additional $\alpha^{t}_{I}$ learning process in APCFR+, which increases its running time. However, we also note that the running time of SAPCFR+ is nearly identical to that of PCFR+, as the only difference between their implementations is a single line of code, which does not alter the computational complexity. Importantly, the computational complexity remains exactly the same, even with no change in the constant factors.
@@ 6 Conclusions
We propose APCFR+, a novel variant of PCFR+. APCFR+ employs the asynchronization of step-sizes in the updates of implicit and explicit accumulated counterfactual regrets to make a more conservative prediction than PCFR+. This conservative prediction reduces the the impact of the prediction inaccuracy, which typically slows down the convergence rate of PCFR+, to improve the robustness. In addition, we introduce SAPCFR+, requiring only a single line modification to PCFR+. Experimental results validate that APCFR+ and SAPCFR+ exhibit a faster empirical convergence rate than PCFR+. To the best of our knowledge, we are the first to propose the asynchronization of step-sizes in the updates of implicit and explicit accumulated counterfactual regrets, a simple yet novel technique that effectively improves the robustness of PCFR+.
Moreover, the techniques used in other PCFR+ variants are compatible with our algorithm. By integrating the techniques from other PCFR+ variants, we introduce APDCFR+ and Stable APDCFR+, which further improve the empirical convergence rate of APCFR+ in HUNL Subgames.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@
[…]
$\small\bm{x}^{t+1}:=\arg\min_{\bm{x}^{\prime}\in\mathcal{D}}\left\{\langle\bm{\ell}^{t},\bm{x}^{\prime}\rangle+\frac{1}{\eta}\mathcal{B}_{\psi}(\bm{x}^{\prime}\parallel\bm{x}^{t})\right\},$ (19)
@@
From the analysis in Section D of Farina et al. (2021), by setting $\psi(\cdot)$ as the the quadratic regularizer $\frac{1}{2}\|\cdot\|^{2}_{2}$, the update rule APCFR+ at infoset $I$ can be written as
@@
where $\eta>0$ is a constant. Note that if $\bm{\theta}^{0}_{I}=\bm{0}$ for all $I\in\mathcal{I}$, then for any $\eta$, the strategy profile sequence $\{\sigma^{1},\sigma^{2},\cdots,\sigma^{T}\}$ generated by APCFR+ is same and $\eta\bm{R}^{t}=\bm{\theta}^{t}$.
@@
[…]
$\small\bm{x}^{t+1}:=\arg\min_{\bm{x}\in\mathcal{D}}\left\{\langle\bm{\ell}^{t},\bm{x}\rangle+\frac{1}{\eta}\mathcal{B}_{\psi}(\bm{x}\parallel\bm{x}^{t})\right\}$
@@
$\small\langle\bm{\ell}^{t},\bm{x}^{t+1}-\bm{x}^{\prime}\rangle\leq\frac{1}{\eta}\left(\mathcal{B}_{\psi}(\bm{x}^{\prime},\bm{x}^{t})-\mathcal{B}_{\psi}(\bm{x}^{\prime},\bm{x}^{t+1})-\mathcal{B}_{\psi}(\bm{x}^{t+1},\bm{x}^{t})\right).$
@@
$\quad\langle-{\bm{r}}^{t}_{I},{\bm{\theta}}^{t+1}_{I}-\sigma_{i}(I)\rangle\leq\frac{1}{\eta}\left(\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t+1}_{I})-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})\right)$ (21)
$\Leftrightarrow$ $\quad\langle-{\bm{r}}^{t}_{I},{\bm{\theta}}^{t+1}_{I}-\hat{\bm{\theta}}^{t}_{I}+\hat{\bm{\theta}}^{t}_{I}-\sigma_{i}(I)\rangle\leq\frac{1}{\eta}\left(\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t+1}_{I})-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})\right).$
@@
$\leq$ $\frac{1}{\eta}\left(\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t+1}_{I})-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})\right)+\frac{1+\alpha^{t}_{I}}{\eta}\left(\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},\hat{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}({\hat{\bm{\theta}}}^{t}_{I},{\bm{\theta}}^{t}_{I})\right),$
@@
$\frac{1+\alpha^{t}_{I}}{\eta}\left(\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},\hat{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\hat{\bm{\theta}}^{t}_{I},{\bm{\theta}}^{t}_{I})\right)$

### v2

@@ 5 Experiments
Figure 3: Empirical convergence rates of the tested algorithms in standard commonly used IIG benchmarks. In all plots, the x-axis is the number of iterations, and the y-axis is exploitability, displayed on a logarithmic scale. Liar’s Dice ($x$) represents that every player is given a die with $x$ sides. Goofspiel ($x$) denotes that each player is dealt $x$ cards. Battleship ($x,y,z$) implies the size of the grid is $x\times y$, and the number of shots is $z$.
Configurations. We now evaluate the empirical convergence rates of APCFR+ and SAPCFR+ by comparing them to PCFR+, Stable PCFR+, Smooth PCFR+, Reg-CFR (liu2022power), and Clairvoyant CFR (farina2023regret). Stable PCFR+ and Smooth PCFR+ are advanced PCFR+ variants. Reg-CFR and Clairvoyant CFR achieve theoretical convergence rates of $O(1/T^{\frac{3}{4}})$ and $O(1/T)$, respectively, while that of other algorithms is $O(1/\sqrt{T})$. Following the settings in PCFR+, we employ alternating updates for both APCFR+ and SAPCFR+. For Stable PCFR+, Smooth PCFR+, and Reg-CFR, we apply alternating updates, as described in their original paper or open-source code. Clairvoyant CFR does not utilize alternating updates, in accordance with its original design. For all algorithms, we utilize quadratic averaging. For all compared algorithms, we adopt the hyperparameters as suggested in their respective original versions. Details on the size of the tested games are in Appendix D (Table 3). The experiments are conducted on a machine equipped with a Xeon(R) Gold 6444Y CPU and 256 GB of memory.
@@
Empirical convergence rates in standard IIG benchmarks. We now present the empirical convergence rates across five standard IIG benchmarks, e.g., Kuhn Poker, Leduc Poker, Goofspiel Poker, Liar’s Dice, and Battleship. These games are implemented using OpenSpiel (lanctot2019openspiel). The algorithm implementations are based on LiteEFG (liu2024liteefgefficientpythonlibrary), as LiteEFG provides approximately 100 times speedup compared to the default implementation in OpenSpiel. The results are in Figure 3. For most of the tested games, except for Battleship (3,2,3) and Goofspiel (4), APCFR+ and SAPCFR+, significantly outperform all baselines. Even in Battleship (3,2,3) and Goofspiel (4), APCFR+ and SAPCFR+ outperform all algorithms except PCFR+. Remarkably, they exhibit performance comparable to PCFR+, reaching similar levels of exploitability after 5000 iterations. Based on the experimental results in Appendix D (Figures 6 and 7), we observe that in the games where our algorithms perform similar to PCFR+, such as Battleship (3,2,3) and Goofspiel (4), PCFR+ also exhibits a rapid decrease in the inaccuracy between the predicted and observed instantaneous counterfactual regrets (detailed discussions are in Appendix D). Furthermore, the performance gap between APCFR+ and SAPCFR+ is relatively small. Specifically, APCFR+ only outperforms SAPCFR+ in Leduc Poker, Battleship (3,2,3), and Liar’s Dice (5). This small performance gap means that in practical applications, SAPCFR+ can be directly used due to its ease of implementation and a faster empirical convergence rate compared to PCFR+. Regarding Stable PCFR+ and Smooth PCFR+, we find that they significantly underperform PCFR+. Reg-CFR and Clairvoyant CFR significantly underperform relative to other algorithms.
Empirical convergence rates in HUNL Subgames. To assess the performance of APCFR+ and SAPCFR+ in addressing real-world games, we also conduct evaluations in HUNL Subgames, which are considerably larger than standard IIG benchmarks. Despite the presence of code related to HUNL Subgames in Openspiel, we have not successfully executed it. Therefore, we utilize HUNL Subgames implemented by Poker RL (steinberger2019pokerrl). More precisely, our code is based on the code from xu2024minimizing. The code in xu2024minimizing supports only Subgame 3 and Subgame 4, so we conduct experiments solely on these two HUNL Subgames. We do not compare Reg-CFR and Clairvoyant CFR in HUNL Subgames, as they perform significantly worse than other CFR algorithms, even in standard IIG benchmarks. The results are shown in Table 1: APCFR+ and SAPCFR+ consistently outperform all baselines in both subgames.
@@
Running times. To validate the efficiency of APCFR+ and SAPCFR+, we compare their running time with that of PCFR+ under the same number of iterations (i.e., 5000). The experimental results are in Appendix D (Table 4). The running time of APCFR+ is slightly higher compared to PCFR+, primarily due to the additional $\alpha^{t}_{I}$ learning process in APCFR+. However, the running time of SAPCFR+ is nearly identical to that of PCFR+, as the only difference between their implementations is a single line of code, which does not alter the computational complexity. Notably, the computational complexity remains exactly the same, even with no change in the constant factors.
PCFR+ Stable PCFR+ Smooth PCFR+ APCFR+ SAPCFR+
Subgame 3 1.44e-3 1.41e-3 (-2.1%) 1.42e-3 (-1.4%) 1.02e-3 (-29.2%) 9.44e-4 (-34.4%)
Subgame 4 1.04e-3 9.77e-4 (-5.3%) 1.02e-3 (-1.9%) 7.53e-4 (-27.6%) 7.83e-4 (-24.7%)
Table 1: Final exploitability for the tested algorithms in HUNL Subgames. Values in red indicate percentages relative to PCFR+.
Leduc Poker (5) Leduc Poker (9) Leduc Poker (13)
DCFR 2.79e-5 1.27e-5 1.09e-5
PCFR+ 2.69e-5 5.21e-5 3.15e-5
APCFR+ 4.80e-6 (-82.1%) 4.03e-5 (-22.6%) 1.45e-5 (-54.0%)
SAPCFR+ 3.49e-6 (-87.0%) 4.07e-5 (-21.9%) 1.42e-5 (-55.0%)
DCFR+ 1.15e-5 (-58.8%) 6.41e-6 (-49.5%) 8.56e-6 (-21.6%)
APDCFR+ 3.69e-6 (-86.3%, -86.7%) 3.42e-6 (-93.4%, -73.1%) 3.02e-6 (-90.4%, -72.3%)
Table 2: The final exploitability for DCFR, PCFR+, APCFR+, SAPCFR+, DCFR+, and APDCFR+ in Leduc Poker variants. Values in red indicate percentages of PCFR+ variants relative to PCFR+, and values in blue indicate percentages of DCFR variants relative to DCFR. Notably, APDCFR+ can serve as a variant of both PCFR+ and DCFR.
Dynamics of $\alpha^{t}_{I}$ in APCFR+. To study the behavior of $\alpha^{t}_{I}$, we analyze its dynamics, as shown in Appendix D (Figures 4 and 5). We observe that $\alpha^{t}_{I}$ experiences a rapid increase during the initial phase but ceases to grow after approximately 100 iterations. This might be due to that the values of $\|{\bm{r}}^{t}_{I}-{\bm{r}}^{t-1}_{I}\|^{2}_{2}$ and $\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}$ are significantly larger in the initial phase than at later stages (as also observed in Figures 6 and 7). More details are in Appendix D.
Dynamics of $\sum_{t=1}^{T}\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$, $\sum_{t=1}^{T}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}$, $\sum_{t=1}^{T}(\frac{\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2})$, and $\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\frac{{\bm{r}}^{t-1}_{I}}{{1+\alpha^{t}_{I}}}\|^{2}_{2}$. To evaluate the regret bound presented in our theoretical analysis, we examine the dynamics of these terms, as demonstrated in Appendix D. Specifically, the dynamics of $\sum_{t=1}^{T}\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$ and $\sum_{t=1}^{T}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}$ are presented in Figures 6 and 7. Similarly, Figures 8 and 9 present the dynamics of $\sum_{t=1}^{T}\left(\frac{\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}\right)$. Additionally, Figures 10 and 11 depict the dynamics of $\sum^{T}_{t=1}\|\bm{r}^{t}_{I}-\frac{\bm{r}^{t-1}_{I}}{1+\alpha^{t}_{I}}\|^{2}_{2}$. This experimental results show that $\|{\bm{r}}^{t}_{I}-{\bm{r}}^{t-1}_{I}\|^{2}_{2}$ are larger than $\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}$, which confirms that APCFR+ effectively reduces the impact of $\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}$ by increasing the weights on $\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}$. For the first three terms, their values are smaller in APCFR+ and SAPCFR+ compared to PCFR+, implying a lower regret bound in Theorem 4.1. However, the value of $\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\frac{{\bm{r}}^{t-1}_{I}}{{1+\alpha^{t}_{I}}}\|^{2}_{2}$ significantly exceeds that of $\sum_{t=1}^{T}(\frac{\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2})$, indicating that the regret bound in Theorem B.1 is extremely higher than that in Theorem 4.1. Thus, we use Theorem 4.1 in the main text instead of Theorem B.1.
Empirical convergence rates of APCFR+ with an alternative learning approach for $\alpha^{t}$. We also experiment with a different learning approach for $\alpha^{t}$, other than the one in Eq. (10), e.g., $\alpha^{t}_{I}=\min\left(\sqrt{\frac{\max_{\tau\in[t-1]}\|\bm{r}^{\tau}_{I}-\bm{r}^{\tau-1}_{I}\|^{2}_{2}}{\max_{\tau\in[t-1]}\|\bm{R}^{\tau+1}_{I}-\bm{R}^{\tau}_{I}\|^{2}_{2}}},\alpha_{max}\right)$, where we also set $\alpha_{max}=5$ as did in Eq. (10) to reduce the cost of hyperparameter tuning. The results in Appendix D (Figure 12 and Table 5) indicate that this approach performs similarly to the one in Eq. (10).
Comparison with other classical CFR algorithms and the generalization of our approach. In addition to the CFR algorithms that have already been compared, we also compare APCFR+ and SAPCFR+ with the classic CFR algorithms: CFR, CFR+, and DCFR. Initially, we conducted experiments using standard IIG benchmarks and HUNL Subgames (Figure 13 and Table 6), where APCFR+ and SAPCFR+ consistently outperformed CFR and CFR+ across all games. However, in poker games like Leduc Poker and HUNL Subgames, APCFR+ and SAPCFR+ did not surpass DCFR. Notably, our algorithms and DCFR are not mutually exclusive and can be combined effectively. The core innovation of our algorithms—the asymmetry of step sizes—can be integrated with DCFR, which involves discounting prior iterations when calculating accumulated regrets. Therefore, we propose APDCFR+ by combining APCFR+ with DCFR (details of APDCFR+ are in Appendix C). In addition to CFR, CFR+, DCFR, APCFR+, and SAPCFR+, we also compare APDCFR+ with DCFR+ (xu2024minimizing), which is an advanced variant of DCFR. Experimental results, detailed in Appendix D (Table 6), demonstrate that APDCFR+ achieves a substantially faster empirical convergence rate compared to the other evaluated algorithms.
Additionally, to further evaluate the performance of DCFR, PCFR+, APCFR+, SAPCFR+, DCFR+, and APDCFR+ in poker games, we conduct tests on various Leduc Poker variants that are used in the original PCFR+ paper (farina2021faster). Specifically, we test on Leduc Poker with ranks of 5, 9, or 13. We denote these Leduc Poker variants as Leduc Poker ($x$), where $x$ represents the number of ranks, noting that the original Leduc Poker has 3 ranks. The results, in Table 2, demonstrate that APCFR+ and SAPCFR+ consistently outperform PCFR+ across all Leduc Poker variants. Notably, the degree to which APCFR+ and SAPCFR+ surpass PCFR+ does not depend on the size of the game. Specifically, the smallest improvement of APCFR+ and SAPCFR+ over PCFR+ occurs in Leduc Poker (9), where the reduction in exploitability is less than half of the reduction observed in Leduc Poker (13). Moreover, the results indicate that DCFR does not consistently outperform PCFR+. For instance, in Leduc Poker (5), the performance of DCFR is inferior to that of PCFR+. More importantly, APDCFR+ consistently outperforms all other algorithms across each Leduc Poker variant tested, except in Leduc Poker (5), where it slightly underperforms compared to SAPCFR+.
@@ 6 Conclusions
We propose a novel variant of PCFR+ called APCFR+, which employs the adaptive asymmetry of step sizes in the updates of implicit and explicit accumulated counterfactual regrets to improve the robustness of PCFR+. We also introduce SAPCFR+, requiring only a single line modification to PCFR+. Experimental results validate that APCFR+ and SAPCFR+ exhibit a faster empirical convergence rate than PCFR+. To our knowledge, we are the first to propose the asymmetry of step sizes in the updates of implicit and explicit accumulated counterfactual regrets, a simple yet novel technique that effectively improves the robustness of PCFR+. Moreover, the techniques used in other CFR+ algorithms are compatible with our algorithm, which shows the generalization of our approach. For example, for DCFR, by using our approach, we propose APDCFR+, which significantly outperforms DCFR in poker games. Future work involves designing more effective $\alpha^{t}$ learning approaches to further enhance the empirical convergence rate.
Acknowledgements
This work is supported in part by the National Natural Science Foundation of China under Grants 62192783 and 62506157, the Jiangsu Science and Technology Major Project BG2024031, the Fundamental Research Funds for the Central Universities (14380128), the Collaborative Innovation Center of Novel Software Technology and Industrialization, and the InnoHK funding.
@@ Acknowledgements
To prove Theorem 4.1, we use the equivalence between RM+ and Online Mirror Descent (OMD) proposed by farina2021faster. We first introduce OMD. OMD is a traditional regret minimization algorithm (nemirovskij1983problem). Let $\bm{\ell}^{t}\in\mathbb{R}^{d}$, $\bm{x}^{t}\in\mathcal{D}$, and let $\psi:\mathcal{D}\to\mathbb{R}^{d}_{\geq 0}=\{\bm{y}|\bm{y}\in\mathbb{R}^{d},\bm{y}\geq\bm{0}\}$ be a 1-strongly convex differentiable regularizer with respect to some norm $\|\cdot\|$, OMD generates the decisions via
$\bm{x}^{t+1}:=\arg\min_{\bm{x}^{\prime}\in\mathcal{D}}\left\{\langle\bm{\ell}^{t},\bm{x}^{\prime}\rangle+\frac{1}{\eta}\mathcal{B}_{\psi}(\bm{x}^{\prime}\parallel\bm{x}^{t})\right\},$ (15)
@@
From the analysis in Section D of farina2021faster, by setting $\psi(\cdot)$ as the quadratic regularizer $\frac{1}{2}\|\cdot\|^{2}_{2}$, the update rule APCFR+ at infoset $I$ can be written as
@@
where $\eta>0$ is a constant and $\sigma^{t}_{i}(I)=\frac{\hat{\bm{\theta}}^{t}_{I}}{\|\hat{\bm{\theta}}^{t}_{I}\|_{1}}$. Note that if $\bm{\theta}^{0}_{I}=\bm{0}$ for all $I\in\mathcal{I}$, then for any $\eta$, the strategy profile sequence $\{\sigma^{1},\sigma^{2},\cdots,\sigma^{T}\}$ generated by APCFR+ is same and $\eta\bm{R}^{t}_{I}=\bm{\theta}^{t}_{I}$ (farina2021faster).
@@
[Lemma 4 of farina2021faster] Let $\mathcal{D}\subseteq\mathbb{R}^{d}$ be closed and convex, let $\bm{\ell}^{t}\in\mathbb{R}^{d}$, $\bm{x}^{t}\in\mathcal{D}$, and let $\psi:\mathcal{D}\to\mathbb{R}_{\geq 0}$ be a 1-strongly convex differentiable regularizer with respect to some norm $\|\cdot\|$. Then,
$\bm{x}^{t+1}\in\arg\min_{\bm{x}\in\mathcal{D}}\left\{\langle\bm{\ell}^{t},\bm{x}\rangle+\frac{1}{\eta}\mathcal{B}_{\psi}(\bm{x}\parallel\bm{x}^{t})\right\}$ (17)
@@
$\langle\bm{\ell}^{t},\bm{x}^{t+1}-\bm{x}^{\prime}\rangle\leq\frac{1}{\eta}\left(\mathcal{B}_{\psi}(\bm{x}^{\prime},\bm{x}^{t})-\mathcal{B}_{\psi}(\bm{x}^{\prime},\bm{x}^{t+1})-\mathcal{B}_{\psi}(\bm{x}^{t+1},\bm{x}^{t})\right).$ (18)
@@
$\langle-{\bm{r}}^{t}_{I},{\bm{\theta}}^{t+1}_{I}-\sigma_{i}(I)\rangle\leq\frac{1}{\eta}\left(\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t+1}_{I})-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})\right)$ (19)
$\Leftrightarrow$ $\langle-{\bm{r}}^{t}_{I},{\bm{\theta}}^{t+1}_{I}-\hat{\bm{\theta}}^{t}_{I}+\hat{\bm{\theta}}^{t}_{I}-\sigma_{i}(I)\rangle\leq\frac{1}{\eta}\left(\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t+1}_{I})-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})\right).$
@@
$\leq$ $\frac{1}{\eta}\left(\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\sigma_{i}(I),{\bm{\theta}}^{t+1}_{I})-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})\right)+$
$\quad\quad\quad\frac{1+\alpha^{t}_{I}}{\eta}\left(\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},\hat{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}({\hat{\bm{\theta}}}^{t}_{I},{\bm{\theta}}^{t}_{I})\right),$
@@
$\quad\quad\quad\frac{1+\alpha^{t}_{I}}{\eta}\left(\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}({\bm{\theta}}^{t+1}_{I},\hat{\bm{\theta}}^{t}_{I})-\mathcal{B}_{\psi}(\hat{\bm{\theta}}^{t}_{I},{\bm{\theta}}^{t}_{I})\right)$